# Cauchy activation function and XNet

### v1

@@ 3.4 Poisson Equation with Dirichelet Boundary Condition
Since the loss values were initially large and decreased significantly over time, the overall loss curve did not clearly highlight the differences between the methods. So we plotted the loss curve for the final 1000 epochs separately, where it becomes evident that our method (Cauchy activation) holds a clear advantage over the original PINN.
Figure 13: Loss curve for PINN and XNet
@@ 3.4.1 High Dimensional PDE
Assume for all $s,t\in[0,T]$, $x,w\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, $z\in\mathbb{R}^{1\times d}$, $m\in\mathbb{N}$ that $\gamma_{m}=5\cdot 10^{-4}$, $d=100$, $T=\frac{3}{10}$, $N=20$, $\mu(t,x)=0$, $\sigma(t,x)w=\sqrt{2}w$, $\xi=(0,0,\ldots,0)\in\mathbb{R}^{d}$, $\Upsilon(s,t,x,w)=x+\sqrt{2}w$, $f(t,x,y,z)=y-y^{3}$, and $g(x)=\left[2+\frac{2}{5}\|x\|_{\mathbb{R}^{d}}^{2}\right]^{-1}$. Note that the solution $u$ satisfies for all $t\in[0,T)$, $x\in\mathbb{R}^{d}$ that $u(T,x)=g(x)$ and
$\frac{\partial u}{\partial t}(t,x)+u(t,x)-[u(t,x)]^{3}+\left(\Delta_{x}u\right)(t,x)=0.$

### v2

@@ 3.5 Poisson Equation with Dirichelet Boundary Condition
Since the loss values were initially large and decreased significantly over time, the overall loss curve did not clearly highlight the differences between the methods. To address this, we plotted the loss curve for the final 1000 epochs separately, making it evident that our method (Cauchy activation) holds a clear advantage over the original PINN.
Figure 20: Overall loss curve for PINN using Tanh and Cauchy activations.
Figure 21: Zoomed-in loss curve for the final 1000 epochs.
Figure 22: Comparison of loss curves for PINN using Tanh and Cauchy activations. The Cauchy activation shows superior convergence speed and stability.
@@
3.5.1 Burger’s equation
To validate the performance of our proposed Cauchy activation function, we included the Burger equation in our experiments. The Burger equation, commonly used in fluid dynamics, is defined as:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},$
where $u(x,t)$ represents the velocity field, and $\nu$ is the viscosity coefficient.
In our implementation, we leveraged a Physics-Informed Neural Network (PINN) to solve the equation. The network approximates the solution $u(x,t)$ and enforces the equation’s physical constraints by minimizing the residual:
$f(x,t)=\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}.$
Modifications for Experiments
We simplified the network architecture by reducing its depth from 10 layers to 5 layers compared to the original design. This adjustment improved computational efficiency while maintaining the capacity to accurately model the solution.
In addition to this, we replaced traditional activation functions, such as tanh, with our custom Cauchy activation function in all layers. This modification was intended to leverage the rapid convergence and robust learning characteristics of the Cauchy activation function.
Results
The results demonstrated a remarkable improvement in training efficiency:
  • With the tanh activation function, the network required 1000 epochs to reduce the training loss to zero.
  • With the Cauchy activation function, the training loss converged to zero within 20 epochs, showcasing the superior convergence speed of our approach.
Figure 23: Training loss comparison for Burger’s equation using tanh and Cauchy activation functions.
These findings validate that the Cauchy activation function is particularly well-suited for physics-informed neural networks, where rapid convergence and efficient training are critical.
@@ 3.5.2 High Dimensional PDE
The Allen-Cahn equation is a reaction-diffusion equation that arises in physics, serving as a prototype for the modeling of phase separation and order-disorder transitions (see, e.g., [10]). This equation is defined as:
$\frac{\partial u}{\partial t}(t,x)+u(t,x)-[u(t,x)]^{3}+\left(\Delta_{x}u\right)(t,x)=0,$
with the solution $u$ satisfying for all $t\in[0,T)$, $x\in\mathbb{R}^{d}$:
$u(T,x)=g(x).$
Assume for all $s,t\in[0,T]$, $x,w\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, $z\in\mathbb{R}^{1\times d}$, $m\in\mathbb{N}$ that $d=100$, $T=\frac{3}{10}$, $\mu(t,x)=0$, $\sigma(t,x)w=\sqrt{2}w$, and $\Upsilon(s,t,x,w)=x+\sqrt{2}w$. The reaction term is defined as
$f(t,x,y,z)=y-y^{3},$
capturing the double-well potential of the Allen-Cahn equation, where the two minima at $y=-1$ and $y=1$ represent stable equilibrium states.
The terminal condition
$g(x)=\left[2+\frac{2}{5}\|x\|_{\mathbb{R}^{d}}^{2}\right]^{-1}$
ensures smooth decay for large $\|x\|$, aligning with the expected physical behavior in bounded domains. These assumptions simplify the Allen-Cahn equation and provide a well-posed high-dimensional problem to evaluate the performance of the proposed method.